# Can Few-shot Work in Long-Context? Recycling the Context to Generate Demonstrations

### v3

@@ Abstract
Despite recent advancements in Large Language Models (LLMs), their performance on tasks involving long contexts remains sub-optimal. In-Context Learning (ICL) with few-shot examples may be an appealing solution to enhance LLM performance in this scenario; However, naïvely adding ICL examples with long context introduces challenges, including substantial token overhead added for each few-shot example and context mismatch between the demonstrations and the target query. In this work, we propose to automatically generate few-shot examples for long context QA tasks by recycling contexts. Specifically, given a long input context (1-3k tokens) and a query, we generate additional query-output pairs from the given context as few-shot examples, while introducing the context only once. This ensures that the demonstrations are leveraging the same context as the target query while only adding a small number of tokens to the prompt. We further enhance each demonstration by instructing the model to explicitly identify the relevant paragraphs before the answer, which improves performance while providing fine-grained attribution to the answer source. We apply our method on multiple LLMs and obtain substantial improvements (+16 absolute points on average across models) on various QA datasets with long context, especially when the answer lies within the middle of the context. Surprisingly, despite introducing only single-hop ICL examples, LLMs also successfully generalize to multi-hop long-context QA using our approach.
@@
Can Few-shot Work in Long-Context?
Recycling the Context to Generate Demonstrations
@@ 1 Introduction
Figure 1: Performance of Gemini Flash (v1.5) on a sample of the Lost-in-the-middle dataset (Liu et al., 2023). The X-axis is the position of the relevant passage in the context. The baseline (blue line) displays a U-shaped curve, performing well only when the relevant passage is at the beginning or end of the input. The oracle (green line) shows significant performance gain when the relevant passage ID is provided in the prompt, showing that the identification of supporting evidence(s) is a major challenge. DoubleDipper (our method, orange line) flattens this U-shaped trend.
Long contexts are prevalent in various domains, ranging from legal documents and scientific articles to lengthy reports and novels. These may consist of a single extensive document or multiple passages, typically retrieved through specific retrieval mechanisms (e.g., RAG; Lewis et al., 2020).
Yet, while Large Language Models (LLMs) have demonstrated impressive capabilities in a variety of tasks including answering questions requiring one or multiple reasoning steps, they often struggle to answer simple questions when faced with long contexts. Despite substantial engineering efforts (Chen et al., 2023) to extend the context window of LLMs to extremely long inputs (32k and even 1M tokens), these models continue to struggle with much shorter inputs, comprising only a few thousand tokens.
In order to answer questions from long inputs, models should implicitly identify relevant information segments and then reason over these segments to formulate an answer. It has been shown that LLMs struggle when the relevant information is buried in the middle of the context (Liu et al., 2023) or obscured by numerous irrelevant details (Levy et al., 2024). Our analysis (illustrated in Figure 1) identifies the identification of relevant information as a major performance bottleneck of current models in long contexts.
In this work, we introduce a novel method to enhance the QA performance of LLMs in long input setups (to allow direct comparisons across a wide swath of models, we limit "long context" here to 1-3k tokens). Our approach, termed DoubleDipper, leverages LLMs’ In-Context Learning (ICL) capability and is based on two principles. First, instead of typical ICL, where each few-shot example is standalone with a separate length context and a question-answer (QA) pair, we propose to recycle the given input context and automatically generate few-shot examples from this context. Specifically, we randomly select a few paragraphs from the given input context and generate QA pairs for each passage. These generated QAs serve as demonstration examples and are placed between the input context and the target input question. Figure 2 illustrates the differences between the traditional ICL with few-shot examples and DoubleDipper. Second, we further enrich each ICL example to instruct the model to explicitly identify the paragraph(s) containing the relevant information before generating the answer. This can be regarded as a structured Chain of Thought that incentivizes the model to pinpoint relevant information before reasoning, an essential capability for long-context processing.
@@
By generating few-shot demonstrations from various sections of the input context while instructing the model to identify relevant passages, DoubleDipper encourages the model to develop deeper reading comprehension skills specific to the given input evidence. This, in turn, allows the model to answer subsequent queries with higher accuracy. DoubleDipper presents several advantages. In terms of efficiency, since each example does not include its own input context, our method adds to the original prompt a minimal number of tokens, resulting in a substantially cheaper inference than traditional ICL. Additionally, recycling the same context for ICL demonstrations ensures that the few-shot examples refer to the same domain as the input question, thus obviating the need for external retrieval processes. Finally, DoubleDipper generates answers with attribution to relevant paragraphs, improving the model’s lookup ability and offering transparency, which substantially simplifies human evaluation (Slobodkin et al., 2024).
@@
We applied DoubleDipper to a variety of LLMs, both commercial (Gemini Pro, Nano and Flash; Reid et al., 2024) and open-source (Llama 3.1 (Dubey et al., 2024), Mistral (Jiang et al., 2023), Mixtral (Jiang et al., 2024) and Gemma (Riviere et al., 2024)), and evaluate it on various QA datasets with long inputs, including common multi-hop QA datasets. Our experiments demonstrate that with only 3 self-generated few-shot examples, DoubleDipper consistently outperforms the baseline on our evaluation set by 16 absolute points on average across models. In addition, for some models, DoubleDipper enhances the robustness to the position of the relevant information within the text. Interestingly, while our few-shot examples focus on single-paragraph answers, DoubleDipper generalizes well to multi-hop QAs and where the answer requires information from multiple passages.
@@ 2 Background
LLMs have been well-documented to struggle when input length grows (An et al., 2023), and especially so when it exceeds input lengths seen during training (Anil et al., 2022). Various methods have been proposed to advance long-context capabilities: Architectural, e.g., to augment the embedding layer to cleverly extrapolate to unseen lengths (Vaswani et al., 2017; Press et al., 2021; Caciularu et al., 2022; Hsieh et al., 2024); and via data, e.g., to incorporate longer inputs and more challenging long-context scenarios into training (He et al., 2023; Chen et al., 2023). However, this challenging problem stubbornly remains in competitive models today (Liu et al., 2023; Bishop et al., 2023; Levy et al., 2024). In contrast to the above methods, DoubleDipper does not involve training or architectural changes.
In documenting and exploring LLM performance in long-context settings, many different benchmarks targeting it have been proposed, such as Scrolls and Zero-Scrolls (Shaham et al., 2022, 2023), Loogle (Li et al., 2023), LongBench (Bai et al., 2023), inter alia. The problem of designing informative and reliable benchmarks in long-context is an an active, ever-changing area of research (Goldman et al., 2024; Yen et al., 2024). We describe the most relevant evaluation benchmarks used in this work in Section 1.
@@
The area of in-context learning (ICL) is a class of prompting techniques where demonstrations are added to the prompt in order to steer or improve model behavior (Min et al., 2022a). Typically, in-context learning involves hand-crafted demonstrations (Song et al., 2022), automatic retrieval of demonstrations from a larger set (Paranjape et al., 2023), or instructing the model to perform various tasks one after another in a pipeline (Gao et al., 2022). Recent improvements in long-context capabilities of LLMs have also had effect on improving the yield from in-context learning by simply using more short-length demonstrations (Agarwal et al., 2024).
While such methods are widely used for their effectiveness (Brown et al., 2020b; an Luo et al., 2024), they remain under-explored in settings of long-context. The reason is simple: If the context is already extremely long, adding additional demonstrations comparable in length to the input context will likely amplify the existing limitations of long context handling (Li et al., 2024c).
More related to our work, a few recent studies propose to prompt LLMs for automatically generating in-context demonstrations for various short context tasks (Kim et al., 2022; Yasunaga et al., 2024; Li et al., 2024a, b). However, applying those methods to long contexts is not straightforward because LLMs would need to generate long demonstrations.
In this work, we tackle these challenges and present a novel ICL method that recycles the given long context and further instructs models to identify the relevant information before generating the answer.
@@ 3 DoubleDipper
Recall that our work focuses on the task of question answering (QA) with long input context comprising multiple paragraphs. In addition to the answer, we aim to identify the supporting paragraphs in order to provide attribution. Formally, given a long input text $D$ composed of $n$ paragraphs $D=\{p_{1},p_{2},...,p_{n}\}$ and a question $q$, the goal is to generate the answer $a$ and identify the set(s) of paragraphs that support the answer $S=\{s_{1},...,s_{k}\}$. The number of the supporting paragraphs is not known in advance and can be one or more.
@@
Figure 3 illustrates DoubleDipper. Starting with the input paragraphs $D$, we initially select $k$ paragraphs at random (e.g., paragraphs $15$, $5$, and $17$, for $k:=3$). For each chosen paragraph, we prompt the model to formulate a question that pertains to the specific paragraph, accompanied by an appropriate answer (for further details on prompt specifications, refer to Appendix A). Each generated QA pair is directly associated with its origin paragraph, enabling us to assemble the following structured in-context demonstration, shown as the DoubleDipper block in Figure 3:
@@
Here, $p_{i}$ indicates the index of the paragraph associated with the QA pair $(q_{i},a_{i})$. Given a test question $q$, we then form a QA prompt by concatenating the original input context $D$, the compiled demonstrations and $q$. The model first generates the one or more indices of the supporting paragraph(s), followed by the answer.
Unlike traditional few shot examples that instruct the model about a specific task, DoubleDipper aims to coach the model on how to “handle” the specific input context. This is achieved by guiding the model to explicitly localize relevant information before generating the answer. Also, by randomly sampling multiple paragraphs from the input, DoubleDipper guarantees that the ICL demonstrations involve reading different parts of the context, allowing the model to better comprehend the input text. Beyond improving the performance of the QA task, instructing the model to provide the supporting paragraphs is valuable on its own as it offers transparency and substantially eases human evaluation (Slobodkin et al., 2024).
DoubleDipper offers several advantages. First, as each example in the demonstration consists only of a question, an answer and the ID of relevant passage, the number of added tokens due to the extra demonstrations is minimal, leading to a low additional cost and computation compared to the traditional In-Context-Learning. Furthermore, by reusing the same context to generate demonstrations, our approach guarantees that all few shot examples are derived from the exact same domain as the input query (Rubin et al., 2022).
Finally, we observe that, although the QA pairs in the demonstration are confined to individual paragraphs, the actual query $q$ may require reasoning over multiple paragraphs (i.e., multi-hop QA). Surprisingly, LLMs can generalize from DoubleDipper local examples to these complex, global questions and successfully generate indices to multiple paragraphs (see Section 5).
@@ 4 Experiments
Size Context Window
Gemini v1.5 Nano 32k
Flash 32k
Pro 32k
Llama 3.1 7B 128k
Mistral v0.3 7B 32k
Mixtral v0.1 7x8B 32k
Gemma 2 2B 8k
9B 8k
Table 1: Models used with DoubleDipper.
@@
Table 2: Evaluation datasets in our experiments. The average number of tokens is computed according to Gemma’s tokenization of the simple prompt.
@@
The Lost-in-the-middle dataset (Liu et al., 2023) includes examples from NaturalQuestions-Open (Kwiatkowski et al., 2019; Lee et al., 2019). Each instance consists of twenty Wikipedia passages, with only one passage containing the answer to the query. The remaining passages are distractors that are lexically similar but do not contain the answer. To assess the robustness of large language models (LLMs) to the position of relevant information, Liu et al. (2023) evaluated cases where the relevant passage appeared in positions 1, 5, 10, 15, and 20. Following their methodology, we sampled 500 instances for each position, resulting in a total of 2,500 instances.
@@
FLenQA (Levy et al., 2024) is a benchmark that includes simple questions with answers of either “True” or “False” based on two key sentences. FLenQA includes three subtasks. The first subtask is Monotone Relations (MonoRel), where each instance asks whether a transitive relation between two entities holds based on the context (e.g., "Is X younger than Y?" based on the sentences "X is younger than Z" and "Z is younger than Y"). The second subtask, People In Rooms (PIR), involves one key sentence indicating that a person is in a specific room and another key sentence describing a property of this room. The question asks whether the person is in a room with the described property. The final subtask is Simplified Rule Taker (SRT), based on RuleTaker (Clark et al., 2020). Each instance consists of a logical rule, two sentences each introducing a fact, and a question over the rule and facts. For each subtask, FLenQA includes contexts with varying lengths, from 50 to 3,000 tokens, by simply adding irrelevant text, demonstrating consistent performance degradation with increased input length. In our experiments, we sampled 250 instances for each subtask with input lengths of 2,000 and 3,000 tokens, leading to a total of 1,500 instances for FLenQA.
@@
In addition, we evaluate our method on common multi-hop QA benchmarks. We sampled 500 instances from each of the following datasets: HotPotQA (Yang et al., 2018), 2Wiki (Ho et al., 2020), and MuSiQue (Trivedi et al., 2021). In all these datasets, the input text includes multiple passages, and models need to perform at least two steps of reasoning over different passages in order to answer the question.

### v4

# DoubleDipper: Improving Long-Context LLMs via Context Recycling

@@ Abstract
Despite recent advancements in Large Language Models (LLMs), their performance on tasks involving long contexts remains sub-optimal. In this work, we propose DoubleDipper, a novel In-Context-Learning method that automatically generates few-shot examples for long context QA tasks by recycling contexts. Specifically, given a long input context (1-3k tokens) and a query, we generate additional query-output pairs from the given context as few-shot examples, while introducing the context only once. This ensures that the demonstrations are leveraging the same context as the target query while only adding a small number of tokens to the prompt. We further enhance each demonstration by instructing the model to explicitly identify the relevant paragraphs before the answer, which improves performance while providing fine-grained attribution to the answer source. We apply our method on multiple LLMs and obtain substantial improvements (+16 absolute points on average across models) on various QA datasets with long context. Surprisingly, despite introducing only single-hop ICL examples, LLMs successfully generalize to multi-hop long-context QA using our approach.
@@
DoubleDipper: Improving Long-Context LLMs via Context Recycling
@@ 1 Introduction
Long contexts are prevalent in various domains, ranging from legal documents and scientific articles to lengthy reports and novels. These may consist of a single extensive document or multiple passages, typically retrieved through specific retrieval mechanisms (e.g., RAG Lewis et al. (2020)).
@@
Yet, while Large Language Models (LLMs) have demonstrated impressive capabilities in a variety of tasks including answering questions requiring one or multiple reasoning steps, they often struggle to answer simple questions when faced with long contexts. Despite substantial engineering efforts (Chen et al., 2023b) to extend the context window of LLMs to extremely long inputs (32k and even 1M tokens), these models continue to struggle with much shorter inputs, comprising only a few thousand tokens, especially when the relevant information is buried in the middle of the context (Liu et al., 2023b) or obscured by numerous irrelevant details (Levy et al., 2024).
In-Context Learning (ICL) with few-shot examples may be an appealing solution to enhance LLM performance in long contexts. However, applying ICL in real-world scenarios without access to training data introduces significant challenges. Developers need to maintain a demonstration pool for retrieving semantically similar demonstrations to any given query (Liu et al., 2022; Rubin et al., 2022). Furthermore, ICL adds a substantial token overhead to the prompt, an issue that becomes even more pronounced with long-context inputs.
In this work, we introduce a novel method to enhance the QA performance of LLMs in long input setups. Our approach, termed DoubleDipper, leverages LLMs’ In-Context Learning capability and is based on two principles. First, instead of typical ICL, where each few-shot example is standalone with a separate lengthy context and a question-answer (QA) pair, we propose to recycle the given input context and automatically generate few-shot examples from this context. Specifically, we randomly select a few paragraphs from the given input context and generate QA pairs for each passage. These generated QAs serve as demonstration examples and are placed between the input context and the target input question. Figure 1 illustrates the differences between the traditional ICL with few-shot examples and DoubleDipper. Second, we enhance each ICL demonstration by explicitly instructing the model to identify relevant information prior to generating an answer. Explicitly identifying relevant passages can be regarded as a structured Chain of Thought that incentivizes the model to pinpoint relevant information before reasoning, an essential capability for long-context processing.
@@
By generating few-shot demonstrations from various sections of the input context while instructing the model to identify relevant passages, DoubleDipper encourages the model to develop deeper reading comprehension skills specific to the given input evidence. This, in turn, allows the model to answer subsequent queries with higher accuracy. DoubleDipper presents several advantages. First, recycling the same context for ICL demonstrations ensures that the few-shot examples refer to the exact same domain as the input question, thus obviating the need for external retrieval of similar demonstrations. Also, in terms of efficiency, since each example does not include its own input context, our method adds to the original prompt a minimal number of tokens, resulting in a substantially cheaper inference than traditional ICL. Finally, DoubleDipper generates answers with attribution to relevant paragraphs, improving the model’s lookup ability and offering transparency, which substantially simplifies human evaluation (Menick et al., 2022; Gao et al., 2023; Liu et al., 2023a; Slobodkin et al., 2024).
@@
We applied DoubleDipper to 12 LLMs, both commercial (Gemini Pro, Nano, Flash (Reid et al., 2024); GPT-4 (Achiam et al., 2023)) and open-source ranging from 2B to 70B parameters (Llama 3.1 (Dubey et al., 2024); Mistral (Jiang et al., 2023); Mixtral (Jiang et al., 2024); Gemma, (Riviere et al., 2024)). We evaluate our method on 7 QA datasets with long inputs, including common multi-hop QA datasets. Our experiments demonstrate that with only 3 self-generated few-shot examples, DoubleDipper consistently outperforms the baseline on our evaluation set by 16 absolute points on average across models. In addition, for some models, DoubleDipper enhances the robustness to the position of the relevant information within the text. Interestingly, while our few-shot examples focus on single-paragraph answers, DoubleDipper generalizes well to multi-hop QAs and where the answer requires information from multiple passages.
@@ 2 Background
LLMs have been well-documented to struggle when input length grows (An et al., 2023), and especially when it exceeds input lengths seen during training (Anil et al., 2022). Various methods have been proposed to advance long-context capabilities: Architectural, e.g., to augment the embedding layer to cleverly extrapolate to unseen lengths (Vaswani et al., 2017; Press et al., 2021; Caciularu et al., 2022; Tan et al., 2024); via data, e.g., to incorporate longer inputs and more challenging long-context scenarios into training (Chen et al., 2024b; He et al., 2024; Chen et al., 2024a); via attention intervention (Hsieh et al., 2024) or by considering question likelihood as a signal for prompt reordering (Liu et al., 2024b). However, this challenging problem stubbornly remains in competitive models (Liu et al., 2023b; Bishop et al., 2024; Levy et al., 2024). In contrast to the above methods, DoubleDipper is a simple method that does not involve training or architectural changes.
Many benchmarks targetting long-context have been proposed, such as Scrolls and Zero-Scrolls (Shaham et al., 2022, 2023), Loogle (Li et al., 2023), LongBench (Bai et al., 2023b), L-Eval (An et al., 2024), inter alia. The problem of designing informative and reliable benchmarks in long-context is an an active, ever-changing area of research (Goldman et al., 2024; Yen et al., 2024). We describe the most relevant evaluation benchmarks used in this work in Section 4.
@@
In-context Learning (ICL) consists of adding demonstrations to the prompt in order to steer or improve model behavior (Min et al., 2022). These demonstrations are either hand-crafted (Song et al., 2022), or retrieved from a large set of training examples (Liu et al., 2022; Rubin et al., 2022; Paranjape et al., 2023). While ICL provides a flexible approach to learn new tasks without updating parameters (Brown et al., 2020b; an Luo et al., 2024), applying ICL in real-world scenarios is challenging notably because there is no available training data for each user query.
More closely related to our work, a few recent studies propose to prompt LLMs for automatically generating in-context demonstrations for various short context tasks. For instance, Kim et al., 2022 focus on sentence classification tasks and prompt LLMs to generate full demonstrations conditioned on a label (e.g., “write a negative review”) and (Chen et al., 2023a; Yasunaga et al., 2024; Li et al., 2024) generate relevant exemplars to the query for reasoning problems. While effective for reasoning tasks with short contexts, these methods are not directly applicable to long-context scenarios because LLMs would need to generate not just a question and an answer, but also an entire long context for each demonstration, which is both computationally expensive and prone to hallucination. DoubleDipper addresses these challenges by introducing a novel and efficient strategy: rather than generating entirely new contexts, it recycles the input context and generates only the question-answer (QA) pairs needed as demonstrations.
@@ 3 DoubleDipper
Given a long input text $\mathcal{C}$ composed of $n$ paragraphs $\mathcal{C}=\{p_{1},p_{2},...,p_{n}\}$ and a question $q$, the goal is to generate the answer $a$ and identify the set(s) of paragraphs that support the answer $S=\{s_{1},...,s_{k}\}$. The number of the supporting paragraphs is not known in advance and can be one or more.
@@
Figure 2 illustrates DoubleDipper. Starting with the input paragraphs $\mathcal{C}$, we initially select $k$ paragraphs at random (e.g., paragraphs 151515, 555, and 171717, for $k:=3$). For each chosen paragraph, we prompt the model to formulate a question that pertains to the specific paragraph, accompanied by an appropriate answer (for further details on prompt specifications, refer to Appendix B). Each generated QA pair is directly associated with its origin paragraph, enabling us to assemble the following structured in-context demonstration, shown as the DoubleDipper block in Figure 2:
@@
Here, $p_{i}$ indicates the index of the paragraph associated with the QA pair $(q_{i},a_{i})$. Given a context $\mathcal{C}$ and a test question $q$, we compile a list of generated demonstrations $\mathcal{D}_{\text{demo}}=(q_{1},p_{1},a_{1},\dots,q_{k},p_{k},a_{k})$ to predict the output $y\sim p_{\theta}(y\mid\mathcal{C},\mathcal{D}_{\text{demo }},q)$ where the output $y$ is the concatenation of one or more indices of the supporting paragraph(s) $S$ and the answer $a$.
Unlike traditional few shot examples that instruct the model about a specific task, DoubleDipper aims to coach the model on how to “handle” the input context. This is achieved by guiding the model to explicitly localize relevant information before generating the answer. Also, by randomly sampling multiple paragraphs from the input, DoubleDipper guarantees that the ICL demonstrations involve reading different parts of the context, allowing the model to better comprehend the input text. Beyond improving the performance of the QA task, instructing the model to provide the supporting paragraphs offers transparency and substantially eases human evaluation.
DoubleDipper offers several advantages. First, as each example in the demonstration consists only of a question, an answer and the ID of relevant passage, the number of added tokens due to the extra demonstrations is minimal (5%), leading to a low additional cost and computation compared to the traditional In-Context-Learning. Furthermore, by reusing the same context to generate demonstrations, our approach guarantees that all few shot examples are derived from the exact same domain as the input query (Rubin et al., 2022).
@@ 4 Experiments
Table 1: Evaluation datasets in our experiments. The average number of tokens is computed according to Gemma’s tokenization of the vanilla prompt.
@@
The Lost-in-the-middle dataset (Liu et al., 2023b) includes examples from NaturalQuestions-Open (Kwiatkowski et al., 2019). Each instance consists of twenty Wikipedia passages, with only one passage containing the answer to the query. The remaining passages are distractors that are lexically similar but do not contain the answer. To assess the robustness of LLMs to the position of relevant information, Liu et al. (2023b) evaluated cases where the relevant passage appeared in positions 1, 5, 10, 15, and 20. Following their methodology, we sampled 500 instances for each position, resulting in a total of 2,500 instances.
@@
FLenQA (Levy et al., 2024) is a benchmark that includes simple questions with answers of either “True” or “False” based on two key sentences. FLenQA includes three subtasks. The first subtask is MonoRel, where each instance asks whether a transitive relation between two entities holds based on the context (e.g., "Is X younger than Y?" based on the sentences "X is younger than Z" and "Z is younger than Y"). The second subtask, PIR, involves one key sentence indicating that a person is in a specific room and another key sentence describing a property of this room. The question asks whether the person is in a room with the described property. The final subtask is SRT, based on RuleTaker (Clark et al., 2020). Each instance consists of a logical rule, two sentences each introducing a fact, and a question over the rule and facts. For each subtask, FLenQA includes contexts with varying lengths, from 50 to 3,000 tokens, by simply adding irrelevant text, demonstrating consistent performance degradation with increased input length. In our experiments, we sampled 250 instances for each subtask with input lengths of 2,000 and 3,000 tokens, leading to a total of 1,500 instances.
@@
In addition, we evaluate our method on common multi-hop QA benchmarks. We sampled 500 instances from HotPotQA (Yang et al., 2018), 2Wiki (Ho et al., 2020), and MuSiQue (Trivedi et al., 2021). In all these datasets, the input text includes multiple passages, and models need to perform at least two steps of reasoning over different passages in order to answer the question.
Models
We apply DoubleDipper to a variety of models, both commercial and open-source. The commercial models include Gemini 1.5 Pro, Gemini 1.5 Flash (Reid et al., 2024) and GPT-4o-mini (Achiam et al., 2023). The open-source models we tested are Llama 3.1 8B, Llama 3.1 70B (Dubey et al., 2024), Gemma 2B (v2), Gemma 9B (v2) and Gemma 27B (Riviere et al., 2024), Mistral-7B-Instruct (v0.2) (Jiang et al., 2023), Mixtral-8x7B-Instruct (v0.1) (Jiang et al., 2024) and Mistral Nemo Instruct 2407111https://mistral.ai/news/mistral-nemo/.
Baselines
We evaluate DoubleDipper against two main baselines. The first is a vanilla baseline that takes as input the entire context $\mathcal{C}$ and the query $q$ and generates only the answer $a$, a common prompting strategy in recent studies on long context (Liu et al., 2023b; Levy et al., 2024). The second baseline, Zero-shot + Evidence Retrieval, prompts the model in a zero-shot manner to first identify relevant passages before generating the answer, following common practices in generating with attribution (Gao et al., 2023; Slobodkin et al., 2024; Fierro et al., 2024).
Evaluation
We evaluate each dataset with the original evaluation metrics. Namely, we report Accuracy for Lost-in-the-middle (Liu et al., 2023b) and FLenQA (Levy et al., 2024), and Token F1 for HotPotQA (Yang et al., 2018), 2Wiki (Ho et al., 2020) and MuSique (Trivedi et al., 2021).
Implementation Details
We randomly select three passages from the input, each containing at least two sentences, and ask the model to generate a single QA pair for each passage. See Section 6 for an analysis of the number of self-generated demonstrations on the performance
5 Results